# Symmetry transformation group arising from the Laplace–Runge–Lenz vector

### v2

@@ 3.1. Symmetry action on constants of motion
The set of all constants of motion of the Kepler problem forms a vector space. Any infinitesimal symmetry $\hat{\bf X}$, acting on phase space, maps this vector space into itself. The mapping is given by the Poisson bracket, as shown by a direct computation.
Proposition 4.
For any constant of motion $C(r^{i},v^{i})$ in the Kepler problem, the associated infinitesimal symmetry generator ${\rm pr}{\bf X}_{C}$ acts on a function $f(r^{i},v^{i})$ on phase space by
${\rm pr}{\bf X}_{C}f=\{f,C\}$ (3.25)
modulo the equations of motion (2.1).
This yields the following result: the only non-zero actions of the infinitesimal symmetries $\hat{\bf X}_{E}$, $\hat{\bf X}_{L^{j}}$, $\hat{\bf X}_{A^{j}}$, $\hat{\bf X}_{\Theta^{j}}$ on the constants of motion $E$, $L^{i}$, $A^{i}$, and $\Theta^{i}$ are given by
${\rm pr}\hat{\bf X}_{L^{j}}(L^{i})=\epsilon^{ij}{}_{k}L^{k},$ (3.26)
${\rm pr}\hat{\bf X}_{A^{j}}(A^{i})=-2E\epsilon^{ij}{}_{k}L^{k},$ (3.27)
${\rm pr}\hat{\bf X}_{L^{j}}(A^{i})=-{\rm pr}\hat{\bf X}_{A^{i}}(L^{j})=\epsilon^{ij}{}_{k}A^{k},$ (3.28)
${\rm pr}\hat{\bf X}_{L^{j}}(\Theta^{i})=-{\rm pr}\hat{\bf X}_{\Theta^{i}}(L^{j})=\epsilon^{ij}{}_{k}\Theta^{k},$ (3.29)
${\rm pr}\hat{\bf X}_{A^{j}}(\Theta^{i})=-{\rm pr}\hat{\bf X}_{\Theta^{i}}(A^{j})=2|A|^{-1}E(\epsilon^{jik}-\epsilon^{jlk}\Theta_{l}\Theta^{i})L_{k}.$ (3.30)
Additionally, Proposition 4 applied to the brackets (3.13) shows that the 7-dimensional vector space spanned by $E$, $L^{i}$, $M^{i}$ carries a homomorphic representation of the LRL symmetry algebra (3.14). The same statement holds for the symmetry algebra (3.24) and the 7-dimensional vector space spanned by $E$, $L^{i}$, $\Theta^{i}$.
3.2. Dynamical symmetry algebra
Some further aspects of the correspondence between the algebra of constants of motion and the Lie algebra of infinitesimal symmetries will now be discussed.
First, quadratic invariants and Casimirs will be considered. For a given Lie algebra of Poisson brackets on phase space, a quadratic scalar $I$ is invariant if it Poisson commutes with all of the constants of motion $C$ in the Lie algebra. Replacement of the constants of motion in $I$ by their corresponding symmetry generators yields a Casimir element in the enveloping algebra of all symmetry generators.
For the Lie algebra (3.14) generated by $\hat{\bf X}_{E}$, $\hat{\bf X}_{L^{j}}$, $\hat{\bf X}_{M^{j}}$, it is well known that [13] the quadratic invariants consist of (linear combinations of)
$E^{2},\quad M^{j}M_{j}-{\rm sgn}(E)\,L^{j}L_{j}$ (3.31)
which can be verified by a simple computation via the Poisson brackets (3.13).
For the Lie algebra generated by $\hat{\bf X}_{E}$, $\hat{\bf X}_{L^{j}}$, $\hat{\bf X}_{\Theta^{j}}$, there is just the single quadratic invariant $E^{2}$, as seen from the only non-vanishing Poisson bracket (3.24).
Any invariant is manifestly a constant of motion in the Kepler problem. More generally, any function of the 7 basic constants of motion $E$, $L^{i}$, $A^{i}$ is again a constant of motion. This means that the set of all constants of motion is infinite dimensional. As a consequence, the set of all infinitesimal symmetry generators, arising from Noether’s theorem in reverse, forms an infinite-dimensional Lie algebra. In this Lie algebra, the infinitesimal generators of point symmetries comprise a finite dimensional Lie subalgebra, which is generated by $\hat{\bf X}_{E}$ and $\hat{\bf X}_{L^{j}}$, namely time-translations and rotations. Quotienting out this subalgebra leaves an infinite-dimensional Lie algebra comprised of genuine infinitesimal dynamical symmetries.
Thus, the finite-dimensional Lie algebras generated respectively by $\hat{\bf X}_{E}$, $\hat{\bf X}_{L^{j}}$, $\hat{\bf X}_{M^{j}}$, and by $\hat{\bf X}_{E}$, $\hat{\bf X}_{L^{j}}$, $\hat{\bf X}_{\Theta^{j}}$, are particular subalgebras in the infinite-dimensional Lie algebra of genuine infinitesimal dynamical symmetries in the Kepler problem.
@@ 4. Symmetry group generated by the directional LRL vector
where $\hat{\bf X}_{\Theta^{j}}$ is the generator (2.44) in phase space, and $\tau^{j}$ is an arbitrary function of $t$, $\vec{r}$}, $\vec{v}$}. In component form,
${\bf Y}_{\Theta^{j}}(t)$ $=\tau^{j},$ (4.2a)
${\bf Y}_{\Theta^{j}}(r^{i})$ $=\tau^{j}v^{i}+2v^{i}r^{j}-v^{j}r^{i}-\delta^{ij}v^{k}r_{k}+2|A|^{-3}E\epsilon^{ikl}A^{j}L_{l}r_{k},$ (4.2b)
${\bf Y}_{\Theta^{j}}(v^{i})$ $=-\kappa|r|^{-3}\tau^{j}r^{i}+v^{i}v^{j}-\kappa|r|^{-3}r^{i}r^{j}-\delta^{ij}(v^{k}v_{k}-\kappa|r|^{-1})+2|A|^{-3}E\epsilon^{ikl}A^{j}L_{l}v_{k}.$ (4.2c)
Now $\tau^{j}$ can be chosen so that one component of this generator is zero, which will consequently be invariant under the infinitesimal symmetry (4.1).
It will be useful to have $|r|$ be invariant, ${\bf Y}_{\Theta^{j}}(|r|)=0$. This condition is easily computed via
${\bf Y}_{\Theta^{j}}(|r|^{2})$ $=2|r|{\bf Y}_{\Theta^{j}}(|r|)$ (4.3)
$=2r_{i}{\bf Y}_{\Theta^{j}}(r^{i})=(\tau^{j}+r^{j})v^{i}r_{i}-v^{j}r^{i}r_{i},$
which yields
$\tau^{j}=(v^{j}r^{i}-r^{j}v^{i})r_{i}=-\epsilon^{jik}r_{i}L_{k}.$ (4.4)
Hence, this fixes all of the gauge freedom.
The components of the gauge-fixed generator are given by
${\bf Y}_{\Theta^{j}}(t)$ $=-\epsilon^{jik}L_{k}r_{i},$ (4.5a)
${\bf Y}_{\Theta^{j}}(r^{i})$ $=-\epsilon^{jlk}r_{l}L_{k}v^{i}+2v^{i}r^{j}-r^{i}v^{j}-\delta^{ij}v^{k}r_{k}+2|A|^{-3}E\epsilon^{ikl}A^{j}L_{l}r_{k},$ (4.5b)
${\bf Y}_{\Theta^{j}}(v^{i})$ $=\kappa|r|^{-3}\epsilon^{jlk}r_{l}L_{k}r^{i}+v^{i}v^{j}-\kappa|r|^{-3}r^{i}r^{j}-\delta^{ij}(v^{k}v_{k}-\kappa|r|^{-1})+2|A|^{-3}E\epsilon^{ikl}A^{j}L_{l}v_{k}.$ (4.5c)
This generator represents a vector field in the coordinate space $(t,r^{i},v^{i})$, and its flow thereby defines the dynamical symmetry transformations
@@
acting in this space, where $\varepsilon^{j}$ is the parameter of the flow.
The flow (4.6) satisfies the system of 7 coupled differential equations
@@
The goal is to integrate this system explicitly. Hereafter, for notational simplicity, the dependence on $\varepsilon^{j}$ will be suppressed in the transformed variables and the generator.
As a starting point, the chosen gauge condition implies that
$\frac{d{|r|^{*}}}{d{\varepsilon_{j}}}={\bf Y}_{\Theta^{j}}^{*}(|r|^{*})=0,$ (4.9)
and therefore,
$|r|^{*}=|r|.$ (4.10)
This constitutes the solution of one combination of the 7 differential equations in the system (4.7).
The strategy for integrating the remainder of the system will involve the following three steps.
(1) The first step will be to convert the differential equations for $r^{i}{}^{*}$ and $v^{i}{}^{*}$ into an equivalent simpler system with the constants of motion $E{}^{*}$, $L^{i}{}^{*}$, $\Theta^{i}{}^{*}$ taken as the variables in the flow, together with the invariant $|r|$. This system is obtained from the action of the infinitesimal symmetry $\hat{\bf X}_{\Theta^{i}}$ on $E{}^{*}$, $L^{i}{}^{*}$, $\Theta^{i}{}^{*}$, which is given by the Poisson brackets.
(2) The second step will consist of expressing $r^{i}{}^{*}$ and $v^{i}{}^{*}$ algebraically in terms of $E{}^{*}$, $L^{i}{}^{*}$, $\Theta^{i}{}^{*}$, and $|r|$. This amounts to a straightforward linear algebra computation.
(3) The third step will be to directly integrate the remaining differential equation for $t{}^{*}$, as its righthand side will involve only solved quantities.
Details are carried out in the appendix.
@@
$t{}^{*}=t+\int_{0}^{\vec{\varepsilon}}\big(\alpha^{r}{}^{*}\vec{L}^{*}\times\vec{\Theta}+\beta^{r}{}^{*}|\vec{L}^{*}|^{2}\vec{\Theta}\big)\cdot d\vec{\varepsilon},$ (4.11a)
${\vec{r}}{}^{*}=\alpha^{r}{}^{*}\vec{\Theta}+\beta^{r}{}^{*}\vec{L}^{*}\times\vec{\Theta},\quad{\vec{v}}{}^{*}=\alpha^{v}{}^{*}\vec{\Theta}+\beta^{v}{}^{*}\vec{L}^{*}\times\vec{\Theta}$ (4.11b)
@@
$\alpha^{r}{}^{*}=|\vec{L}^{*}|^{2}-\kappa|{\vec{r}}|,\quad\beta^{r}{}^{*}={\rm sgn}({\vec{r}}\cdot{\vec{v}})|{\vec{r}}|\sqrt{2(E+\kappa|{\vec{r}}|^{-1})-(|{\vec{r}}|^{-1}|\vec{L}{\mathstrut}^{*}|)^{2}},$ (4.12a)
$\alpha^{v}{}^{*}=-{\rm sgn}({\vec{r}}\cdot{\vec{v}})\kappa\sqrt{2(E+\kappa|{\vec{r}}|^{-1})-(|{\vec{r}}|^{-1}|\vec{L}{\mathstrut}^{*}|)^{2}},\quad\beta^{v}{}^{*}=2E+\kappa|{\vec{r}}|^{-1},$ (4.12b)
@@
$\vec{L}^{*}=\vec{L}+\vec{\varepsilon}\times\vec{\Theta}.$ (4.13)
Under these transformations, $E$ and $\vec{\Theta}$ are invariant, along with $|{\vec{r}}|$.
As shown by Theorem 1, the three infinitesimal symmetries $\hat{\bf X}_{\Theta^{j}}$ form an abelian Lie algebra $\mathbb{R}^{3}$. This implies that the set of all dynamical symmetry transformations (4.11) under composition generates a 3-dimensional abelian Lie group.
The physical meaning of this symmetry group is that it preserves energy and the direction of the LRL vector, while angular momentum is shifted by a vector in the plane orthogonal to the LRL vector. Hence, the magnitude of the LRL vector is changed:
$|\vec{A}^{*}|=\sqrt{|\vec{A}|^{2}+2E\big(\vec{\varepsilon}\cdot(\vec{\Theta}\times\vec{L})+|\vec{\varepsilon}|^{2}-(\vec{\varepsilon}\cdot\vec{\Theta})^{2}\big)}.$ (4.14)
This change is proportional to the change in eccentricity. In the case of elliptical orbits, since energy is unchanged, the period $T=2\pi/\sqrt{-2E}$ and semi-major axis $a=\kappa/(-2E)$ are also preserved.
@@ 5. LRL symmetry transformation group
where $\hat{\bf X}_{A^{j}}$ is the symmetry generator (2.37) in phase space, and $\tau^{j}$ is an arbitrary function of $t$, $\vec{r}$}, $\vec{v}$}. This gauge freedom will be fixed by taking $|r|$ to be invariant:
${\bf Y}_{A^{j}}(|r|^{2})$ $=2|r|{\bf Y}_{A^{j}}(|r|)$ (5.2)
$=2r_{i}{\bf Y}_{A^{j}}(r^{i})=(\tau^{j}+r^{j})v^{i}r_{i}-v^{j}r^{i}r_{i},$
which yields expression (4.4) for $\tau^{j}$. The gauge-fixed generator thus has the components
${\bf Y}_{A^{j}}(t)$ $=-\epsilon^{jik}L_{k}r_{i},$ (5.3a)
${\bf Y}_{A^{j}}(r^{i})$ $=-\epsilon^{jlk}r_{l}L_{k}v^{i}+2v^{i}r^{j}-r^{i}v^{j}-\delta^{ij}v^{k}r_{k},$ (5.3b)
${\bf Y}_{A^{j}}(v^{i})$ $=\kappa|r|^{-3}\epsilon^{jlk}r_{l}L_{k}r^{i}+v^{i}v^{j}-\kappa|r|^{-3}r^{i}r^{j}-\delta^{ij}(v^{k}v_{k}-\kappa|r|^{-1}),$ (5.3c)
which define a vector field in the coordinate space $(t,r^{i},v^{i})$.
@@
where $\varepsilon^{j}$ is the parameter of the flow.
In component form, the flow is defined by a system of 7 coupled differential equations
@@
To integrate this system, the same three steps used in the previous section will be employed with one change: In the first step, the primary variables will be taken to be $E^{*}$ and $|r|$, together with $M^{i}{}^{*}$ and $L^{i}{}^{*}$, as the latter satisfy a coupled system shown by their Poisson brackets (3.13), with $\Theta^{i}{}^{*}$ no longer being invariant in the flow. This leads to a more complicated transformation on $r^{i}{}^{*}$ and $v^{i}{}^{*}$ in the second step. Details of all three steps are carried out in the appendix.
@@
$\underline{R}(\phi,\hat{n})=\cos(\phi)\ \underline{I}+(1-\cos(\phi))\underline{I}\,\hat{n}\,\hat{n}\cdot+\sin(\phi)\underline{S}\,\hat{\varepsilon}\times$ (5.6)
@@
$\underline{I}=\begin{pmatrix}1&0\\ 0&1\end{pmatrix},\quad\underline{S}=\sqrt{-{\rm sgn}(E)}\begin{pmatrix}0&1\\ {\rm sgn}(E)&0\end{pmatrix}$ (5.7)
@@
$t{}^{*}=t+\int_{0}^{\vec{\varepsilon}}\big(\alpha^{r}{}^{*}\vec{L}^{*}\times\vec{\Theta}^{*}+\beta^{r}{}^{*}|\vec{L}^{*}|^{2}\vec{\Theta}^{*}\big)\cdot d\vec{\varepsilon},$ (5.8a)
${\vec{r}}{}^{*}=\alpha^{r}{}^{*}\vec{\Theta}^{*}+\beta^{r}{}^{*}\vec{L}^{*}\times\vec{\Theta}^{*},\quad{\vec{v}}{}^{*}=\alpha^{v}{}^{*}\vec{\Theta}^{*}+\beta^{v}{}^{*}\vec{L}^{*}\times\vec{\Theta}^{*}$ (5.8b)
with parameter $\vec{\varepsilon}$, where the coefficients are given by expressions (4.12).
@@
in terms of $\vec{M}=|\vec{M}|\vec{\Theta}$ with $|\vec{M}|^{2}={\rm sgn}(E)|\vec{L}|^{2}+\tfrac{1}{2}\kappa^{2}|E|^{-1}$. Under these transformations, $E$ and $|\vec{M}|^{2}-{\rm sgn}(E)\,|\vec{L}|^{2}$, along with $|{\vec{r}}|$, are invariant.
@@
$\vec{\Theta}{}^{*}=\vec{\Theta},\quad\vec{L}^{*}=\vec{L}+|\vec{A}|\,\vec{\varepsilon}\times\vec{\Theta}$ (5.10)
@@
are the transformed constants of motion, with $|A|^{2}=\kappa^{2}+2E|\vec{L}|^{2}$. Under these transformations, $E$, $\vec{\Theta}$, and $|{\vec{r}}|$ are invariant, while the magnitude of the LRL vector is changed:
$|\vec{A}^{*}|=|\vec{A}|\sqrt{1+2E|\vec{A}|{\mathstrut}^{-1}\big(\vec{\varepsilon}\cdot(\vec{\Theta}\times\vec{L})+|\vec{\varepsilon}|^{2}-(\vec{\varepsilon}\cdot\vec{\Theta})^{2}\big)}.$ (5.11)
@@
From Proposition 3, note that the three infinitesimal symmetries $\hat{\bf X}_{A^{j}}$ do not comprise a closed algebra, but instead they generate a 6-dimensional Lie algebra that contains the spatial rotations $\mathfrak{so}(3)$ as a subalgebra generated by $\hat{\bf X}_{L^{j}}$ through the commutator of $\hat{\bf X}_{A^{j}},\hat{\bf X}_{A^{k}}]$. Consequently, the LRL dynamical symmetry transformations (5.8) do not themselves form a group. However, if the parameter $\hat{\varepsilon}$ is taken to be fixed, then the set of resulting transformations generates a one-dimensional Lie group with parameter $\phi=\sqrt{-2E}\,|\vec{\varepsilon}|$. For $E<0$, the group is isomorphic to the rotation group $SO(2)$, while for $E>0$, it is isomorphic to the hyperbolic rotation group $SO(1,1)$. When $E=0$, the group is isomorphic to $\mathbb{R}$. In each case, the composition of all these one-dimensional groups will generate a 6-dimensional Lie group whose structure depends on $E$, which is $SO(4)$ when $E<0$; $SO(3,1)$ when $E<0$; $SO(3)\rtimes\mathbb{R}^{3}$ when $E=0$.

### v3

@@ 3.1. Symmetry action on constants of motion
The set of all constants of motion of the Kepler problem forms a vector space, and any infinitesimal symmetry $\hat{\bf X}$ maps this vector space into itself. The mapping is explicitly given by the Poisson bracket ${\rm pr}{\bf X}_{C}f=\{f,C\}$, as shown by a direct computation. Therefore, the only non-zero actions of the infinitesimal symmetries $\hat{\bf X}_{E}$, $\hat{\bf X}_{L^{j}}$, $\hat{\bf X}_{A^{j}}$, $\hat{\bf X}_{\Theta^{j}}$ on the constants of motion $E$, $L^{i}$, $A^{i}$, and $\Theta^{i}$ are given by
${\rm pr}\hat{\bf X}_{L^{j}}(L^{i})=\epsilon^{ij}{}_{k}L^{k},\quad{\rm pr}\hat{\bf X}_{A^{j}}(A^{i})=-2E\epsilon^{ij}{}_{k}L^{k},\quad{\rm pr}\hat{\bf X}_{L^{j}}(A^{i})=-{\rm pr}\hat{\bf X}_{A^{i}}(L^{j})=\epsilon^{ij}{}_{k}A^{k},$ (3.9a)
${\rm pr}\hat{\bf X}_{L^{j}}(\Theta^{i})=-{\rm pr}\hat{\bf X}_{\Theta^{i}}(L^{j})=\epsilon^{ij}{}_{k}\Theta^{k},\ {\rm pr}\hat{\bf X}_{A^{j}}(\Theta^{i})=-{\rm pr}\hat{\bf X}_{\Theta^{i}}(A^{j})=2|A|^{-1}E\epsilon^{jlk}(\delta_{l}^{i}-\Theta_{l}\Theta^{i})L_{k}.$ (3.9b)
Similarly, from the brackets (3.5), observe that the 7-dimensional vector space spanned by $E$, $L^{i}$, $M^{i}$ carries a homomorphic representation of the LRL symmetry algebra. The same statement holds for the symmetry algebra in Theorem 1 on the span of $E$, $L^{i}$, $\Theta^{i}$.
@@ 4. Symmetry group generated by the directional LRL vector
where $\hat{\bf X}_{\Theta^{j}}$ is the generator (2.30) in phase space, and $\tau^{j}$ is an arbitrary function of $t$, $\vec{r}$}, $\vec{v}$}. Now $\tau^{j}$ can be chosen so that one component of this generator (4.1) is zero, which will consequently be invariant under the infinitesimal symmetry (4.1).
It will be useful to have $r$ be invariant, ${\bf Y}_{\Theta^{j}}(r)=0$. This condition is easily computed via
$r{\bf Y}_{\Theta^{j}}(r)=r_{i}{\bf Y}_{\Theta^{j}}(r^{i})$ $=\tau^{j}r_{i}v^{i}+|A|^{-2}(|L|^{2}-\kappa r)\epsilon^{jkl}\Theta_{k}L_{l}$ (4.2)
which determines
$\tau^{j}=|A|^{-2}(\kappa r-|L|^{2})(r_{i}v^{i})^{-1}\epsilon^{jkl}\Theta_{k}L_{l}.$ (4.3)
The components of the gauge-fixed generator represent a vector field in the coordinate space $(t,r^{i},v^{i})$. Its flow thereby defines the dynamical symmetry transformations
@@
acting in this space, where $\varepsilon^{j}$ is the parameter of the flow. The components of this flow (4.4) satisfy the system of 7 coupled differential equations
@@
The goal now will be to integrate this system explicitly. Hereafter, for notational simplicity, the dependence on $\varepsilon^{j}$ will be suppressed in the transformed variables and the generator.
As a starting point, the gauge condition implies $\frac{d{r^{*}}}{d{\varepsilon_{j}}}={\bf Y}_{\Theta^{j}}^{*}(r^{*})=0$, and therefore
$r^{*}=r.$ (4.7)
constitutes the solution of one combination of the 7 differential equations in the system (4.5). The strategy for integrating the remainder of the system involves three steps: (1) Convert the differential equations for $r^{i}{}^{*}$ and $v^{i}{}^{*}$ into an equivalent simpler system with the constants of motion $E{}^{*}$, $L^{i}{}^{*}$, $\Theta^{i}{}^{*}$ taken as the variables in the flow, together with the invariant $r$. This system is obtained from the action of the infinitesimal symmetry $\hat{\bf X}_{\Theta^{i}}$ on $E{}^{*}$, $L^{i}{}^{*}$, $\Theta^{i}{}^{*}$, which is given by the Poisson brackets. (2) Express $r^{i}{}^{*}$ and $v^{i}{}^{*}$ algebraically in terms of $E{}^{*}$, $L^{i}{}^{*}$, $\Theta^{i}{}^{*}$, and $r$. This amounts to a straightforward linear algebra computation. (3) Directly integrate the remaining differential equation for $t{}^{*}$, as its righthand side will involve only solved quantities. Details of these steps will be given elsewhere [17].
@@
$t{}^{*}=t+\mathcal{T}(r,|\vec{L}^{*}|,E)-\mathcal{T}(r,|\vec{L}|,E),$ (4.8a)
${\vec{r}}{}^{*}=|\vec{A}^{*}|^{-1}(\alpha^{r}{}^{*}\vec{\Theta}+\beta^{r}{}^{*}\vec{L}^{*}\times\vec{\Theta}),\quad{\vec{v}}{}^{*}=|\vec{A}^{*}|^{-1}(\alpha^{v}{}^{*}\vec{\Theta}+\beta^{v}{}^{*}\vec{L}^{*}\times\vec{\Theta})$ (4.8b)
@@
$\vec{L}^{*}=\vec{L}+\vec{\varepsilon}\times\vec{\Theta}$ (4.9)
@@
$\alpha^{r}{}^{*}=|\vec{L}^{*}|^{2}-\kappa|{\vec{r}}|,\quad\beta^{r}{}^{*}=\gamma^{*},\quad\alpha^{v}{}^{*}=-\kappa|{\vec{r}}|^{-1}\gamma^{*},\quad\beta^{v}{}^{*}=2E+\kappa|{\vec{r}}|^{-1},$ (4.10)
with
$\mathcal{T}(r,|\vec{L}|,E)=\begin{cases}\dfrac{1}{2E}\Big(\gamma-\dfrac{\kappa}{\sqrt{2|E|}}\arctan\Big(\dfrac{\sqrt{2|E|}\gamma}{2E|{\vec{r}}|+\kappa}\Big)\Big),&E<0\\ \dfrac{1}{2E}\Big(\gamma-\dfrac{\kappa}{\sqrt{2E}}{\rm arctanh}\Big(\dfrac{\sqrt{2E}\gamma}{2E|{\vec{r}}|+\kappa}\Big)\Big),&E>0\\ \dfrac{\kappa|{\vec{r}}|+|\vec{L}|^{2}}{3\kappa^{2}}\gamma,&E=0\end{cases}$ (4.11)
and
$\gamma={\rm sgn}({\vec{r}}\cdot{\vec{v}})\sqrt{2(E|{\vec{r}}|+\kappa)|{\vec{r}}|-|\vec{L}|^{2}}.$ (4.12)
Under these transformations, $E$ and $\vec{\Theta}$ are invariant, along with $|{\vec{r}}|$, while
$|\vec{A}^{*}|=\sqrt{\kappa^{2}+2E|\vec{L}{}^{*}|}=\sqrt{|\vec{A}|^{2}+2E\big(\vec{\varepsilon}\cdot(\vec{\Theta}\times\vec{L})+|\vec{\varepsilon}|^{2}-(\vec{\varepsilon}\cdot\vec{\Theta})^{2}\big)}.$ (4.13)
As shown by Theorem 1, the three infinitesimal symmetries $\hat{\bf X}_{\Theta^{j}}$ form an abelian Lie algebra $\mathbb{R}^{3}$. This implies that the set of all dynamical symmetry transformations (4.8) under composition generates a 3-dimensional abelian Lie group. Its physical meaning is that the group transformations preserve energy and the direction of the LRL vector, while angular momentum is shifted by a vector in the plane orthogonal to the LRL vector. The change in the magnitude of the LRL vector is proportional to the change in eccentricity. In the case of elliptical orbits, since energy $E<0$ is unchanged, the period $T=2\pi/\sqrt{2|E|}$ and semi-major axis $a=\kappa/(2|E|)$ are also preserved.
@@ 5. LRL symmetry transformation group
where $\hat{\bf X}_{A^{j}}$ is the symmetry generator (2.24) in phase space, and $\tau^{j}$ is an arbitrary function of $t$, $\vec{r}$}, $\vec{v}$}. This gauge freedom will be fixed by taking $r$ to be invariant:
$r{\bf Y}_{A^{j}}(r)=r_{i}{\bf Y}_{A^{j}}(r^{i})=(\tau^{j}+r^{j})v^{i}r_{i}-v^{j}r^{2}$ (5.2)
which yields
$\tau^{j}=r^{2}(v^{i}r_{i})^{-1}v^{j}-r^{j}=(v^{i}r_{i})^{-1}\epsilon^{jkl}L_{k}r_{l}.$ (5.3)
The gauge-fixed generator (5.1) constitutes a vector field in the coordinate space $(t,r^{i},v^{i})$.
@@
where $\varepsilon^{j}$ is the parameter of the flow. In component form, the flow is defined by a system of 7 coupled differential equations
@@
To integrate this system, the same three steps used in the previous section can be employed with one change: In the first step, the primary variables will be taken to be $E^{*}$ and $r$, together with $M^{i}{}^{*}$ and $L^{i}{}^{*}$, as the latter satisfy a coupled system shown by their Poisson brackets (3.5), with $\Theta^{i}{}^{*}$ no longer being invariant in the flow. This leads to a more complicated transformation on $r^{i}{}^{*}$ and $v^{i}{}^{*}$ in the second step. Details will be given elsewhere [17].
@@
$\underline{R}(\phi,\hat{n})=\cos(\phi)\ \underline{I}+(1-\cos(\phi))\underline{I}\,\hat{n}\,\hat{n}\cdot+\sin(\phi)\underline{S}\,\hat{n}\times$ (5.6)
@@
$\underline{I}=\begin{pmatrix}1&0\\ 0&1\end{pmatrix},\quad\underline{S}=\sqrt{-{\rm sgn}(E)}\begin{pmatrix}0&1\\ -{\rm sgn}(E)&0\end{pmatrix}$ (5.7)
@@
$t{}^{*}=t+\int_{0}^{\vec{\varepsilon}}\big(|\vec{L}^{*}|^{2}\vec{\Theta}^{*}+(\kappa|{\vec{r}}|-|\vec{L}^{*}|^{2})\gamma^{*}{}^{-1}\vec{L}^{*}\times\vec{\Theta}^{*}\big)\cdot d\vec{\varepsilon},$ (5.8a)
${\vec{r}}{}^{*}=|\vec{A}^{*}|^{-1}(\alpha^{r}{}^{*}\vec{\Theta}^{*}+\beta^{r}{}^{*}\vec{L}^{*}\times\vec{\Theta}^{*}),\quad{\vec{v}}{}^{*}=|\vec{A}^{*}|^{-1}(\alpha^{v}{}^{*}\vec{\Theta}^{*}+\beta^{v}{}^{*}\vec{L}^{*}\times\vec{\Theta}^{*})$ (5.8b)
with parameter $\vec{\varepsilon}$, where the coefficients $\alpha^{r}$, $\alpha^{v}$, $\beta^{r}$, $\beta^{v}$ are given by expressions (4.10)–(4.12).
@@
in terms of $\vec{M}=|\vec{M}|\vec{\Theta}$ with $|\vec{M}|^{2}={\rm sgn}(E)|\vec{L}|^{2}+\tfrac{1}{2}\kappa^{2}|E|^{-1}$. Under these transformations, $E$ and $|\vec{M}|^{2}-{\rm sgn}(E)\,|\vec{L}|^{2}$, along with $|{\vec{r}}|$, are invariant, while $|\vec{A}^{*}|=\sqrt{\kappa^{2}+2E|\vec{L}{}^{*}|^{2}}$ where
$|\vec{L}^{*}|^{2}$ $=\cos(\sqrt{-2E}\,|\vec{\varepsilon}|)^{2}|\vec{L}|^{2}+\sin(\sqrt{-2E}\,|\vec{\varepsilon}|)^{2}((\hat{\varepsilon}\cdot\vec{L})^{2}+{\rm sgn}(E)(\hat{\varepsilon}\cdot\vec{M})^{2}-{\rm sgn}(E)|\vec{M}|^{2})$ (5.10)
$\qquad+\sqrt{-{\rm sgn}(E)}\sin(2\sqrt{-2E})\vec{\varepsilon}\cdot(\vec{M}\times\vec{L})$
@@
$\vec{\Theta}{}^{*}=\vec{\Theta},\quad\vec{L}^{*}=\vec{L}+\kappa\,\vec{\varepsilon}\times\vec{\Theta}$ (5.11)
@@
are the transformed constants of motion. Under these transformations, $E$, $\vec{\Theta}$, $|\vec{A}|=\kappa$, and $|{\vec{r}}|$ are invariant.
@@
Recall that the three infinitesimal symmetries $\hat{\bf X}_{A^{j}}$ do not comprise a closed algebra when $E\neq 0$, but instead they generate a 6-dimensional Lie algebra that contains the spatial rotations $\mathfrak{so}(3)$ as a subalgebra generated by $\hat{\bf X}_{L^{j}}$ through the commutator of $\hat{\bf X}_{A^{j}},\hat{\bf X}_{A^{k}}]$. Consequently, the LRL dynamical symmetry transformations (5.8) do not themselves form a group. However, if the parameter $\hat{n}=\hat{\varepsilon}$ is taken to be fixed, then the set of resulting transformations generates a one-dimensional Lie group with parameter $\phi=\sqrt{-2E}\,|\vec{\varepsilon}|\neq 0$. For $E<0$, this group is isomorphic to the rotation group $SO(2)$ and acts on $(\vec{L},\vec{M})$ via the planar rotation operator twisted with the Pauli matrix $\underline{\sigma}_{1}$:
$\underline{R}(\sqrt{2|E|},\hat{n})=\cos(\sqrt{2|E|})\ \underline{I}+(1-\cos(\sqrt{2|E|}))\underline{I}\,\hat{n}\,\hat{n}\cdot+\sin(\sqrt{2|E|})\underline{\sigma}_{1}\,\hat{n}\times$ (5.12)
For $E>0$, the group is isomorphic to the hyperbolic rotation group $SO(1,1)$, with its action on $(\vec{L},\vec{M})$ given by the planar boost operator twisted with the Pauli matrix $\underline{\sigma}_{2}$:
$\underline{R}(i\sqrt{2E},\hat{n})=\cosh(\sqrt{2E})\ \underline{I}+(1-\cosh(\sqrt{2E}))\underline{I}\,\hat{n}\,\hat{n}\cdot-\sinh(\sqrt{2E})\underline{\sigma}_{2}\,\hat{n}\times$ (5.13)
In each case, the composition of all these one-dimensional groups with $\hat{n}=\hat{\varepsilon}\in S^{2}$ will generate a 6-dimensional Lie group whose structure depends on $E$, which is $SO(4)$ when $E<0$ and $SO(3,1)$ when $E<0$.